OpenStudy (anonymous):

Zacharias has a total of 65 coins, all of which are either dimes or nickels. The total value of the coins is $5.95. Find the number of each type of coin. Zacharias has a total of 65 coins, all of which are either dimes or nickels. The total value of the coins is $5.95. Find the number of each type of coin. @Mathematics

OpenStudy (zarkon):

D+N=65 .1D+.05N=5.95 solve for D and N

OpenStudy (anonymous):

D + N = 65 0.1D + 0.05N = 5.95

OpenStudy (anonymous):

have to tell solve for D and N